# msbm10 scaled1095

### v1

@@ Abstract.
While the spectral properties and invariant subspaces of compressed shifts on model spaces are well understood, their behaviour on nearly $S^{*}$-invariant subspaces, a natural generalization with weaker structural constraints, remains largely unexplored. In this paper, we completely characterize the point spectrum, whole spectrum and invariant subspace structure for such compressed shifts by unitary equivalence, using the Frostman shift, Crofoot transform, and Sz.-Nagy–Foias theory. Our results reveal how the relaxation of $S^{*}$-invariance impacts spectral structure and invariant subspaces, bridging a gap between classical model space theory and broader function-theoretic settings.
@@ 1. Preliminaries
The study of invariant subspaces for linear operators has led to profound results, bridging complex function theory, operator theory, and functional analysis. A cornerstone in this field is the Beurling theorem, which characterizes shift-invariant subspaces of the Hardy space $H^{2}$ (see, e.g. [18, A. Corollary 1.4.1]). To extend these ideas to Hardy spaces of an annulus, Hitt introduced the concept of nearly $S^{*}$-invariant subspaces in [14], generalizing Hayashi’s work on the kernels of Toeplitz operators [13]. Further contributions by Sarason [19, 20] deepened the connections with Toeplitz kernels, further enriching the theory.
A closed subspace $\mathcal{M}\subseteq H^{2}$ is called nearly $S^{*}$-invariant (originally weakly invariant) if, for any $f\in\mathcal{M}$ with $f(0)=0$, it follows that $S^{*}f\in\mathcal{M}$. Intuitively, $\mathcal{M}$ is nearly $S^{*}$-invariant if the zeros of its functions can be factored out without leaving the space. Hitt’s seminal work provides a complete characterization of such subspaces in $H^{2}$, later extended to a vectorial setting in [4, Theorem 4.4]. More generally it was extended to the case of nearly $S^{*}$-invariant subspaces with defect $m$ (that is, if $f(0)=0$ then $S^{*}f\in\mathcal{M}+\mathcal{F}$ for a fixed subspace $\mathcal{F}$ of minimal dimension $m$) in [5].
@@
[14, Proposition 3] The nearly $S^{*}$-invariant subspaces of $H^{2}$ have the form $\mathcal{M}=hK$, with $h\in\mathcal{M}$ of unit norm, $h(0)>0,$ $h$ orthogonal to all elements of $\mathcal{M}$ vanishing at the origin, $K$ an $S^{*}$-invariant subspace, and the operator of multiplication by $h$ isometric from $K$ into $H^{2}$.
The function $h$ is called the extremal function for $\mathcal{M}$ due to its being the unique solution to the extremal problem
@@
As shown in [11, Proposition 9.14], the invariant subspaces of $S_{\theta}$ in $K_{\theta}$ are precisely $K_{\theta}\ominus K_{\phi}$, where $\phi$ divides $\theta$. This classical result motivates us to ask a natural question as below.
@@
We completely answer Question 1 in Theorem 4.5 and characterize the spectrum of this compressed shift in Theorem 3.9, with both results illustrated in Example 4.6.
@@
with the domain $\mbox{Dom}(A_{g}^{\theta})=\{f\in K_{\theta}:\;gf\in L^{2}\}.$ These operators share many characteristics with classical Toeplitz operators. However, they also exhibit striking differences (see, e.g. [21]). For example, there are bounded truncated Toeplitz operators with unbounded symbols (though any truncated Toeplitz operator with a bounded symbol is itself bounded). Especially, [21, Theorem 3.1] implies $A_{g}^{\theta}=0$ if and only if $g\in\theta H^{2}+\overline{\theta H^{2}}.$
Given a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ with extremal function $h$ and associated inner function $\theta$, $\theta(0)=0$, [12, Lemma 2.4] identified the orthogonal projection $P_{\mathcal{M}}$ from $L^{2}$ onto $\mathcal{M}$ as
@@
As a corollary, [12, Corollary 2.5] exhibits the reproducing kernel for $\mathcal{M}=hK_{\theta}$ as given by
@@
where Dom$(A_{g}^{\mathcal{M}})=\{f\in\mathcal{M}:\;gf\in L^{2}\}.$
The multiplier $M_{h}:K_{\theta}\rightarrow hK_{\theta}$ is a surjective isometry as in Theorem 1.1, and so it is unitary, with adjoint $M_{h}^{*}=M_{h^{-1}},$ and so there follows the result of [12, Theorem 3.1].
@@
We note that Theorem 1.2 and [21, Theorem 3.1] provide a characterization of zero truncated Toeplitz operators on $\mathcal{M}=hK_{\theta}$.
@@
Kernel spaces are crucial for understanding the structure of linear operators, especially Toeplitz operators and truncated Toeplitz operators (see, e.g. [1, 21]). [16, Corollary 3.2] (with $n=1$) states that the kernel of a truncated Toeplitz operator on model space is nearly $S^{*}$-invariant with defect at most $1$. This result can be extended to the operator $A^{\mathcal{M}}_{g}$ acting on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
@@
Suppose that $f\in\ker A^{\mathcal{M}}_{g}$ and $f(0)=0$. We have $f=hk$ with some $k\in K_{\theta}$ and we note that by near $S^{*}$-invariance, it follows that $h(0)\neq 0$. Otherwise every function in $\mathcal{M}$ would vanish at 00, which yields =$\mathcal{M}=\{0\}$. So it follows that $k(0)=0.$ From $A^{\mathcal{M}}_{g}f=0$, Theorem 1.2 implies
@@
and so $k\in\ker A^{\theta}_{|h|^{2}g}$. Hence [16, Corollary 3.2] asserts that there is a one-dimensional subspace $\mathcal{F}=\langle x\rangle$, independent of $k$, such that $S^{*}k=\overline{z}k\in\ker A^{\theta}_{|h|^{2}g}+\mathcal{F}$. Then there is a $\lambda\in\mathbb{C}$ such that $A^{\theta}_{|h|^{2}g}(\overline{z}k-\lambda x)=0.$ We can express this as
@@
Finally, we demonstrate that the truncated Toeplitz operator $A_{g}^{\mathcal{M}}$ on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ are $D$-symmetric with respect to the conjugation $D:=M_{h}CM_{h}^{*}$, where
@@
This paper is structured as follows: First, we derive explicit representations for the truncated Toeplitz operator $A_{|h|^{2}z}^{\theta}$ and its adjoint $A_{|h|^{2}\bar{z}}^{\theta}$ on the model space $K_{\theta}$ with $\theta(0)=0$ in Propositions 2.3 and 2.1, establishing fundamental results on their cyclic vectors and actions on kernel functions in Propositions 2.4 and 2.5. Next, we apply unitary equivalence, using the Frostman shift, Crofoot transform, and Sz.-Nagy–Foias theory to rigorously analyze the spectral properties of $A_{z}^{\mathcal{M}}$, identifying its essential spectrum in Corollary 3.2, point spectrum in Theorem 3.8 and whole spectrum in Theorem 3.9. Finally, we focus on finding the generalized eigenvectors for any eigenvalue of $A_{|h|^{2}\bar{z}}^{\theta}$ with multiplicity $n\geq 2$ to display more finite-dimensional invariant subspaces in Proposition 4.3, and then classify all invariant subspaces for the compressed shift $A_{z}^{\mathcal{M}}$ acting on an arbitrary nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ in Theorem 4.5. Especially, we demonstrate Theorems 3.9 and 4.5 in Example 4.6, thereby unifying these results within a generalized framework of model space theory.
@@ 2. The compressed shifts on nearly invariant subspaces
So we have established the following proposition.
@@
Under the first decomposition in (2.1), (2.6) yields $B_{v}$ as a rank-1 perturbation of $S_{\theta}^{*}$ on $K_{\theta}.$ Moreover, the constraints of the extremal function $h\in\mathcal{M}=hK_{\theta}$ ensure that $|v|\leq 1$.
@@
The only eigenvalue is 00 and the invariant subspaces are model subspaces of $K_{\theta}$, i.e., $\{0\},K_{z},K_{z^{2}},K_{z^{3}}$ and $K_{z^{4}}$. On the other hand, the matrix of $B_{v}$ is
$[B_{v}]=\left(\begin{array}[]{cccc}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \alpha&0&0&0\end{array}\right)$
where $\alpha=\langle|h|^{2},z^{4}\rangle$. Note that $[B_{v}]^{4}=\alpha I$ and there are four distinct eigenvalues, the fourth roots of $\alpha$ (in the case $\alpha\neq 0$). So the lattice of invariant subspaces is rather different, being the linear spans of eigenvectors. Note that the two lattices of invariant subspaces are very different, even though the matrices differ in only one entry.
@@
$A_{v}f=\left\{\begin{array}[]{ll}zf,&f\in K_{\theta\overline{z}},\\ v,&f\in\theta\overline{z}.\end{array}\right.$ (2.9)
Equation (2.9) indicates how $A_{v}$ is a rank-1 perturbation of $S_{\theta}$ on $K_{\theta}.$ For $S_{\theta}$ on $K_{\theta}$, [11, Proposition 9.13] shows that it is a cyclic operator with cyclic vector $k_{0}=1-\overline{\theta(0)}\theta.$ The next proposition focuses on the cyclic vectors of $A_{v}$ and $B_{v}$.
@@
which shows $\bigvee\{u,B_{v}u\}$ contains $S^{*}u$. By induction, $\bigvee\{B_{v}^{n}u:n\geq 0\}$ contains $\bigvee\{S^{*n}u:n\geq 0\}$. Since $\bigvee\{S^{*n}\theta:\;n\geq 1\}=K_{\theta}$ (see, e.g. [11, Proposition 5.15]) and $\theta(0)=0,$ we conclude that
@@
$\bigvee\{A_{v}^{n}1:\;n\geq 0\}$ =$=\bigvee\{(CB_{v}C)^{n}1:\;n\geq 0\}$
=$=\bigvee\{CB_{v}^{n}C1:\;n\geq 0\}$
@@
$\bigvee\{A_{v}^{n}u:\;n\geq 0\}$ =$=\bigvee\{(CB_{v}C)^{n}u:\;n\geq 0\}$
=$=\bigvee\{CB_{v}^{n}Cu:\;n\geq 0\}$
@@ 3. Spectra of the compressed shifts on nearly invariant subspaces
[…]
@@
[…]
@@
In addition, a linear bounded operator $T$ is said to be Fredholm if its range $R(T)$ is closed and both $R(T)^{\perp}$ and the null space $N(T)$ are finite-dimensional. The essential spectrum is defined as $\sigma_{e}(T)=\{\lambda\in\mathbb{C}:\;T-\lambda I\;\mbox{is not Fredholm}\}$, which is invariant under perturbation by compact operators (see [22]). In the case of $S_{\theta}$, its essential spectrum is given by $\sigma_{e}(S_{\theta})=\sigma(\theta)\cap\mathbb{T}$, where $\mathbb{T}$ denotes the unit circle (see e.g. [11, Proposition 9.26] or [21, Lemma 2.5]). Then Proposition 2.3 implies $\sigma_{e}(A_{v})=\sigma(\theta)\cap\mathbb{T}.$ This result, combined with (1.2), leads to the following corollary on the essential spectrum of $A_{z}^{\mathcal{M}}$.
@@
Next, the eigenvalues together with eigenvectors and then the whole spectrum are explored for $A_{v}$ in (2.9) on $K_{\theta}$ with $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. For the case $v=0$, $A_{v}=S_{\theta}$ on $K_{\theta}.$ And for the case $|v|=1$, $A_{v}$ is the same as the unitary operator $U_{v}$ in Clark’s paper [7] with $\theta(0)=0.$ Since the $v=0$ case is well-understood and our subsequent proofs remain valid for this case, we shall henceforth assume $|v|<1$ without loss of generality. At this time, $A_{v}$ is a c.n.u. contraction on a model space $K_{\theta}$ with $\theta(0)=0.$ In consequence, we apply the Sz.-Nagy–Foias theory of characteristic functions for contractions on Hilbert space to delve our deeper exploration.
@@
$\mathfrak{D}_{T}$ and $\mathfrak{D}_{T^{*}}$ are the defect spaces of $T$. Let us first proceed with the concrete analysis on the defect spaces of $A_{v}$ with $|v|<1$.
@@
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$, $B_{v}=A_{v}^{*}$. Then $\mathfrak{D}_{A_{v}}=\mathbb{C}\theta/z,\;\mathfrak{D}_{B_{v}}=\mathbb{C}$.
@@
Thus we conclude that
$\sqrt{I-A_{v}B_{v}}=\sqrt{1-|v|^{2}}P_{K_{z}}.$
Furthermore, using $CA_{v}C=B_{v}$ and (3.1), it yields that
@@
which implies that
$\sqrt{I-B_{v}A_{v}}=\sqrt{1-|v|^{2}}P_{\theta/zK_{z}}.$
Hence the two defect spaces are $\mathfrak{D}_{A_{v}}=\mathbb{C}\theta/z$ and $\mathfrak{D}_{B_{v}}=\mathbb{C}$ with $\mathop{\rm dim}\nolimits\mathfrak{D}_{A_{v}}=\mathop{\rm dim}\nolimits\mathfrak{D}_{B_{v}}=1.$
Considering [23, Theorem 2.3] in the context of our c.n.u. operator $A_{v}$ with $|v|<1$, we observe that $A_{v}$ is a special case of $Z(A)$ in [10, Definition 2.1] with $A\xi=v\xi$ for $\xi\in N=\mathbb{C}$. And then we may use [10, Theorem 3.2] and Lemma 3.3 to deduce that $A_{v}$ ($|v|<1$) on $K_{\theta}$ with $\theta(0)=0$ has the characteristic function
@@
Applying [23, Theorem 2.3] to our operator $A_{v}$ with $|v|<1$, it yields that $A_{v}$ is unitarily equivalent to the canonical compressed shift $S$ on the model space $H^{2}(\mathbb{C})\ominus\Theta_{A_{v}}(\lambda)H^{2}(\mathbb{C}\theta/z)$, where $H^{2}(\mathbb{C})=H^{2}(\mathbb{D}).$ From (3.2), we can write
$\Theta_{A_{v}}(\lambda)H^{2}(\mathbb{C}\theta/z)$ =$=\left\{\Theta_{A_{v}}(\lambda)\left(\frac{\theta}{z}f(z)\right):\;f\in H^{2}(\mathbb{C})\right\}$
=$=\left\{\frac{\theta(\lambda)-v}{1-\overline{v}\theta(\lambda)}f(z):\;f\in H^{2}(\mathbb{C})\right\}$
@@
Therefore, we obtain the unitary model for $A_{v}$ with $|v|<1$ on $K_{\theta}$ with $\theta(0)=0$ as below.

### v2

@@ Abstract.
While the spectral properties and invariant subspaces of compressed shifts on model spaces are well understood, their behaviour on nearly $S^{*}$-invariant subspaces, a natural generalization under weaker structural constraints, remains largely unexplored. These operators are closely related to the Clark-type unitary operators, yet differ from them in several ways. In this paper, we completely characterize the point spectrum, whole spectrum and invariant subspace structure for such compressed shifts by unitary equivalence, using the Frostman shift, Crofoot transform, and Sz.-Nagy–Foias theory. The results reveal how the relaxation of $S^{*}$-invariance impacts spectral structure and invariant subspaces, bridging a gap between classical model space theory and broader function-theoretic settings.
@@ 1. Preliminaries
The study of invariant subspaces for linear operators has led to profound results, bridging complex function theory, operator theory, and functional analysis. A cornerstone in this field is the Beurling theorem, which characterizes shift-invariant subspaces of the Hardy space $H^{2}$ (see, e.g. [21, A. Corollary 1.4.1]). To extend these ideas to Hardy spaces of an annulus, Hitt introduced the concept of nearly $S^{*}$-invariant subspaces in [15], generalizing Hayashi’s work on the kernels of Toeplitz operators [14]. Further contributions by Sarason [22, 23] deepened the connections with Toeplitz kernels, further enriching the theory.
A closed subspace $\mathcal{M}\subseteq H^{2}$ is called nearly $S^{*}$-invariant (originally weakly invariant) if, for any $f\in\mathcal{M}$ with $f(0)=0$, it follows that $S^{*}f\in\mathcal{M}$. Intuitively, $\mathcal{M}$ is nearly $S^{*}$-invariant if the zeros of its functions can be factored out without leaving the space. For the broader classes of operators, we introduce the notion of nearly invariant subspaces, first for a left invertible operator (see [16]) and later expand to general $C_{0}$-semigroups $\{T(t)\}_{t\geq 0}$ on a separable infinite-dimensional Hilbert space (see [17]). In the nearly $S^{*}$-invariant setting, Hitt’s seminal work provides a complete characterization of such subspaces in $H^{2}$ as below.
@@
[15, Proposition 3] The nearly $S^{*}$-invariant subspaces of $H^{2}$ have the form $\mathcal{M}=hK$, with $h\in\mathcal{M}$ of unit norm, $h(0)>0,$ $h$ orthogonal to all elements of $\mathcal{M}$ vanishing at the origin, $K$ an $S^{*}$-invariant subspace, and the operator of multiplication by $h$ isometric from $K$ into $H^{2}$.
Later this is generalized to a vectorial setting in [4, Theorem 4.4]. More generally, it is further extended to nearly $S^{*}$-invariant subspaces with defect $m$ (i.e., if $f(0)=0$ then $S^{*}f\in\mathcal{M}+\mathcal{F}$ for a fixed subspace $\mathcal{F}$ of minimal dimension $m$) in [5], where the closely related concept of almost-invariant subspace is also investigated, along with its vectorial version in [6].
The function $h$ in Hitt’s theorem is called the extremal function for $\mathcal{M}$ due to its being the unique solution to the extremal problem
@@
As shown in [12, Proposition 9.14], the invariant subspaces of $S_{\theta}$ in $K_{\theta}$ are precisely $K_{\theta}\ominus K_{\phi}$, where $\phi$ divides $\theta$. This classical result motivates us to ask a natural question as below.
@@
We completely answer Question 1 in Theorem 4.5 and characterize the spectrum of this compressed shift in Theorem 3.9; both results are illustrated in Example 4.6.
@@
with the domain $\mbox{Dom}(A_{g}^{\theta})=\{f\in K_{\theta}:\;gf\in L^{2}\}.$ These operators share many characteristics with classical Toeplitz operators. However, they also exhibit striking differences (see, e.g. [24]). For example, there are bounded truncated Toeplitz operators with unbounded symbols (though any truncated Toeplitz operator with a bounded symbol is itself bounded). Especially, [24, Theorem 3.1] implies $A_{g}^{\theta}=0$ if and only if $g\in\theta H^{2}+\overline{\theta H^{2}}.$
Given a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ with extremal function $h$ and associated inner function $\theta$, $\theta(0)=0$, [13, Lemma 2.5] identified the orthogonal projection $P_{\mathcal{M}}$ from $L^{2}$ onto $\mathcal{M}$ as
@@
As a corollary, [13, Corollary 2.6] exhibits the reproducing kernel for $\mathcal{M}=hK_{\theta}$ as given by
@@
[…]
The multiplier $M_{h}:K_{\theta}\rightarrow hK_{\theta}$ is a surjective isometry as in Theorem 1.1, and so it is unitary, with adjoint $M_{h}^{*}=M_{h^{-1}},$ and so there follows the result of [13, Theorem 3.1].
@@
We note that Theorem 1.2 and [24, Theorem 3.1] provide a characterization of zero truncated Toeplitz operators on $\mathcal{M}=hK_{\theta}$.
@@
Kernel spaces are crucial for understanding the structure of linear operators, especially Toeplitz operators and truncated Toeplitz operators (see, e.g. [1, 24]). [19, Corollary 3.2] (with $n=1$) states that the kernel of a truncated Toeplitz operator on model space is nearly $S^{*}$-invariant with defect at most $1$. This result can be extended to the operator $A^{\mathcal{M}}_{g}$ acting on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
@@
Suppose that $f\in\ker A^{\mathcal{M}}_{g}$ and $f(0)=0$. We have $f=hk$ with some $k\in K_{\theta}$ and we note that by near $S^{*}$-invariance, it follows that $h(0)\neq 0$. Otherwise every function in $\mathcal{M}$ would vanish at 0, which yields $\mathcal{M}=\{0\}$. So it follows that $k(0)=0.$ From $A^{\mathcal{M}}_{g}f=0$, Theorem 1.2 implies
@@
and so $k\in\ker A^{\theta}_{|h|^{2}g}$. Hence [19, Corollary 3.2] asserts that there is a one-dimensional subspace $\mathcal{F}=\langle x\rangle$, independent of $k$, such that $S^{*}k=\overline{z}k\in\ker A^{\theta}_{|h|^{2}g}+\mathcal{F}$. Then there is a $\lambda\in\mathbb{C}$ such that $A^{\theta}_{|h|^{2}g}(\overline{z}k-\lambda x)=0.$ We can express this as
@@
Finally, we demonstrate that the truncated Toeplitz operator $A_{g}^{\mathcal{M}}$ on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ is $D$-symmetric with respect to the conjugation $D:=M_{h}CM_{h}^{*}$, where
@@
Driven by Question 1, we investigate both the spectral properties and the invariant subspace problem for the compressed shift operator arising in (1.2). This operator is unitarily equivalent to the truncated Toeplitz operator $A_{v}:=A_{|h|^{2}z}^{\theta}$, a rank-1 perturbation of $S_{\theta}$ (see Proposition 2.3). Unlike the Clark unitary operator $U_{\lambda}$ with $|\lambda|=1$ in [8], however, $A_{v}$ exhibits different behavior due to the condition $|v|=|\langle\theta,|h|^{2}\rangle|<1$. So new technical tools and methods are required for characterizing the relevant properties of $A_{v}$.
The construction developed in this paper proceeds as follows. First, we derive explicit representations for the truncated Toeplitz operator $A_{|h|^{2}z}^{\theta}$ and its adjoint $A_{|h|^{2}\bar{z}}^{\theta}$ on the model space $K_{\theta}$ with $\theta(0)=0$ in Propositions 2.3 and 2.1, establishing fundamental results on their cyclic vectors and actions on kernel functions in Propositions 2.4 and 2.5. Next, we apply unitary equivalence, using the Frostman shift, Crofoot transform, and Sz.-Nagy–Foias theory to rigorously analyze the spectral properties of $A_{z}^{\mathcal{M}}$, identifying its essential spectrum in Corollary 3.2, point spectrum in Theorem 3.8 and whole spectrum in Theorem 3.9. Finally, we explore the generalized eigenvectors for any eigenvalue of $A_{|h|^{2}\bar{z}}^{\theta}$ with multiplicity $n\geq 2$ to display more finite-dimensional invariant subspaces in Proposition 4.3, and then classify all invariant subspaces for $A_{z}^{\mathcal{M}}$ acting on an arbitrary nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ in Theorem 4.5. Especially, we demonstrate Theorems 3.9 and 4.5 in Example 4.6, unifying these results within a generalized framework of model space theory.
@@ 2. The compressed shifts on nearly invariant subspaces
Thus we have established the following proposition.
@@
Under the first decomposition in (2.1), (2.6) yields $B_{v}$ as a rank-1 perturbation of $S_{\theta}^{*}$ on $K_{\theta}.$ Moreover, since $1\in K_{\theta}$ with $\theta(0)=0$, it holds $\langle h,h\rangle=\langle 1,1\rangle=1$, so that $\|h\|=1.$ By Cauchy-Schwartz inequality, we deduce
$|v|=\bigl|\langle\theta,|h|^{2}\rangle\bigr|=\bigl|\langle\theta h,h\rangle\bigr|\leq\|\theta h\|\|h\|\leq\|h\|\|h\|=1$
and $|\langle\theta h,h\rangle|=1$ if and only if $\theta h=\beta h$ for some complex number $\beta$. That is $(\theta-\beta)h=0$ which is impossible, so $v=\bigl|\langle\theta,|h|^{2}\rangle\bigr|<1$.
@@
The only eigenvalue is 0 and the invariant subspaces are model subspaces of $K_{\theta}$, i.e., $\{0\},K_{z},K_{z^{2}},K_{z^{3}}$ and $K_{z^{4}}$. On the other hand, the matrix of $B_{v}$ is
$B_{v}]=\left(\begin{array}[]{cccc}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \alpha&0&0&0\end{array}\right)$
where $\alpha=\langle|h|^{2},z^{4}\rangle$. Note that $B_{v}]^{4}=\alpha I$ and there are four distinct eigenvalues, the fourth roots of $\alpha$ (in the case $\alpha\neq 0$). So the lattice of invariant subspaces is rather different, being the linear spans of eigenvectors. Note that the two lattices of invariant subspaces are very different, even though the matrices differ in only one entry.
@@
$A_{v}f=\left\{\begin{array}[]{ll}zf,&f\in K_{\theta\overline{z}},\\ v,&f=\theta\overline{z}.\end{array}\right$. (2.9)
Equation (2.9) indicates how $A_{v}$ is a rank-1 perturbation of $S_{\theta}$ on $K_{\theta}.$ For $S_{\theta}$ on $K_{\theta}$, [12, Proposition 9.13] shows that it is a cyclic operator with cyclic vector $k_{0}=1-\overline{\theta(0)}\theta.$ The next proposition focuses on the cyclic vectors of $A_{v}$ and $B_{v}$.
@@
which shows $\bigvee\{u,B_{v}u\}$ contains $S^{*}u$. By induction, $\bigvee\{B_{v}^{n}u:n\geq 0\}$ contains $\bigvee\{S^{*n}u:n\geq 0\}$. Since $\bigvee\{S^{*n}\theta:\;n\geq 1\}=K_{\theta}$ (see, e.g. [12, Proposition 5.15]) and $\theta(0)=0,$ we conclude that
@@
$\bigvee\{A_{v}^{n}1:\;n\geq 0\}$ $=\bigvee\{(CB_{v}C)^{n}1:\;n\geq 0\}$
$=\bigvee\{CB_{v}^{n}C1:\;n\geq 0\}$
@@
$\bigvee\{A_{v}^{n}u:\;n\geq 0\}$ $=\bigvee\{(CB_{v}C)^{n}u:\;n\geq 0\}$
$=\bigvee\{CB_{v}^{n}Cu:\;n\geq 0\}$
@@ 3. Spectra of the compressed shifts on nearly invariant subspaces
$\theta(z)=e^{iy}\underbrace{z^{n}\left(\prod_{n\geq 1}\frac{a_{n}}{|a_{n}|}\frac{a_{n}-z}{1-\overline{a_{n}}z}\right)}_{B}\underbrace{\exp\left(-\int_{\mathbb{T}}\frac{\xi+z}{\xi-z}d\mu(\xi)\right)}_{s_{\mu}},$
@@
In other words, $\sigma(\theta)$ consists of the complement of the set of all points $\lambda\in\overline{\mathbb{D}}$ for which the function $1/\theta$ can be analytically continued into a neighborhood of $\lambda.$ This definition aligns with the standard conception in analytic function theory, where the spectrum is the set of all “singularities” of the function in question. Additionally, a useful characterization of $\sigma(\theta)$ in terms of $\liminf$ of zero set of $\theta$ is provided in [12, Proposition 7.19], which states
@@
In addition, a linear bounded operator $T$ is said to be Fredholm if its range $R(T)$ is closed and both $R(T)^{\perp}$ and the null space $N(T)$ are finite-dimensional. The essential spectrum is defined as $\sigma_{e}(T)=\{\lambda\in\mathbb{C}:\;T-\lambda I\;\mbox{is not Fredholm}\}$, which is invariant under perturbation by compact operators (see [25]). In the case of $S_{\theta}$, its essential spectrum is given by $\sigma_{e}(S_{\theta})=\sigma(\theta)\cap\mathbb{T}$, where $\mathbb{T}$ denotes the unit circle (see e.g. [12, Proposition 9.26] or [24, Lemma 2.5]). Then Proposition 2.3 implies $\sigma_{e}(A_{v})=\sigma(\theta)\cap\mathbb{T}.$ This result, combined with (1.2), leads to the following corollary on the essential spectrum of $A_{z}^{\mathcal{M}}$.
@@
Next, the eigenvalues together with eigenvectors and then the whole spectrum are explored for $A_{v}$ in (2.9) on $K_{\theta}$ with $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. The case $v=0$ reduces to the classical compressed shift $S_{\theta}$ on $K_{\theta},$ which is well understood. We therefore focus on the new setting $0<|v|<1$. At this time, $A_{v}$ is a c.n.u. contraction on a model space $K_{\theta}$ with $\theta(0)=0.$ In consequence, the Sz.-Nagy–Foias theory of characteristic functions for contractions on Hilbert space becomes an indispensable tool for our subsequent explorations.
@@
$\mathfrak{D}_{T}$ and $\mathfrak{D}_{T^{*}}$ are the defect spaces of $T$. Let us first proceed with the concrete analysis on the defect spaces of $A_{v}$.
@@
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$, $B_{v}=A_{v}^{*}$. Then $\mathfrak{D}_{A_{v}}=\mathbb{C}\theta/z,\;\mathfrak{D}_{B_{v}}=\mathbb{C}$.
@@
Thus we conclude that $\sqrt{I-A_{v}B_{v}}=\sqrt{1-|v|^{2}}P_{K_{z}}.$ Furthermore, using $CA_{v}C=B_{v}$ and (3.1), it yields that
@@
which implies that $\sqrt{I-B_{v}A_{v}}=\sqrt{1-|v|^{2}}P_{\theta/zK_{z}}.$ Hence the two defect spaces are $\mathfrak{D}_{A_{v}}=\mathbb{C}\theta/z$ and $\mathfrak{D}_{B_{v}}=\mathbb{C}$ with $\mathop{\rm dim}\nolimits\mathfrak{D}_{A_{v}}=\mathop{\rm dim}\nolimits\mathfrak{D}_{B_{v}}=1.$
Considering [26, Theorem 2.3] in the context of our c.n.u. operator $A_{v}$, we observe that $A_{v}$ can be viewed as a special case of $Z(A)$ in [11, Definition 2.1] with $A\xi=v\xi$ for $\xi\in N=\mathbb{C}$. Rather surprisingly, we use [11, Theorem 3.2] and Lemma 3.3 to deduce that $A_{v}$ on $K_{\theta}$ with $\theta(0)=0$ has the characteristic function
@@
This yields that $A_{v}$ is unitarily equivalent to the canonical compressed shift $S$ on a new model space $H^{2}(\mathbb{C})\ominus\Theta_{A_{v}}(\lambda)H^{2}(\mathbb{C}\theta/z)$, where $H^{2}(\mathbb{C})=H^{2}(\mathbb{D}).$ From (3.2), we can write
$\Theta_{A_{v}}(\lambda)H^{2}(\mathbb{C}\theta/z)$ $=\left\{\Theta_{A_{v}}(\lambda)\left(\frac{\theta}{z}f(z)\right):\;f\in H^{2}(\mathbb{C})\right\}$
$=\left\{\frac{\theta(\lambda)-v}{1-\overline{v}\theta(\lambda)}f(z):\;f\in H^{2}(\mathbb{C})\right\}$
@@
Therefore, we obtain the unitary model for $A_{v}$ on $K_{\theta}$ with $\theta(0)=0$ as below.